OpenStudy (anonymous):

solve by the addition method 2x+y=11 x-y=11

OpenStudy (anonymous):

add both the equations

OpenStudy (anonymous):

it will be 3x = 22 so x = 22/3 then put the value of x in any of the above equation to get the value of y

OpenStudy (anonymous):

but whats the value of y??

OpenStudy (anonymous):

13/3??

OpenStudy (anonymous):

from equation (ii) x - y = 11 so y = x - 11 the value of x is 22/3 so the value of y will be 22/3 - 11

OpenStudy (anonymous):

y = -11/3

OpenStudy (anonymous):

ok thank you can u help me to solve by graphing?? only this question?? 3x+4y=20 3x-2y=8

OpenStudy (anonymous):

okay, but you need to know how to graph straight lines to solve using graphing

OpenStudy (anonymous):

do you know how to graph straight lines???

OpenStudy (anonymous):

yes but i just need the points thats all

OpenStudy (anonymous):

yes, then the point where the two lines intersect is your required answer

OpenStudy (anonymous):

yes but what are these points??

OpenStudy (anonymous):

okay, i'll tell you from the very start. do you have graph paper right now???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ill make one

OpenStudy (anonymous):

do you have to solve this specific question using graph???

OpenStudy (anonymous):

no i only need the points the answer :(

OpenStudy (anonymous):

yes, right, i think :P lol

OpenStudy (anonymous):

please tell me the points!!